65=((71*80.4)+(10*31.6)+(x*21))/(71+10+x)
solve for x

Answer by checkley77(12844)

65=((71*80.4)+(10*31.6)+(x*21))/(71+10+x)
65=(5,708.4+316+21x)/(81+x)
65=(6,024.4+21x)/(81+x)
65(81+x)=6,024.4+21x
5,265+65x=6,024.4+21x
65x-21x=6,024.4-5,265
44x=759.4
x=759.4/44
x=17.25909 ans.